We roll 2 dice and sum the numbers on the faces.

a) What is the probability that we roll a 5?

b) What is the probability that we roll a 2 or a 3?

c) What is the probability that we roll at least a 10?

d) What is the probability that we roll at most a 5?

e) What is the probability that we roll an even sum?

f) What is the probability that we roll either an even sum, or a sum of at least 9?
Answer #1

Number of outcomes when 2 dices are rolled = 36

(a) 5 = (2,3), (3,2), (1,4), (4,1)i. e. 4 ways

Required probability = 4/36 = 1/9​​​​​​

(b) 2 or 3 = (1,1)(1,2),(2,1) i. e 3 ways

Required probability = 3/36 = 1/12​​​​​​

(c) atleast 10 = (5,5),(4,6),(6,4),(5,6),(6,5),(6,6) i. e 6 ways

Required probability = 6/36 = 1/6

(d) atmost 5 = (1,1),(1,2),(2,1),(2,2),(1,3),(3,1)(2,3),(3,2),(1,4),(4,1) i. e 10 ways

Required probability = 10/36 = 5/18

(e) even sum (2,4,6,8,10,12) = (1,1), (2,2),(1,3),(3,1),(1,5),(5,1)(2,4),(4,2),(3,3)(3,5)(5,3),(4,4)(2,6),(6,2)(5,5)(4,6),(6,4)(6,6) i. e 18 ways

Required probability = 18/36 = 1/2​​​​​​

(f) even sum or sum of atleast 9 = all the outcomes of part (e) just include (9,11) = (4,5),(5,4)(3,6),(6,3),(5,6),(6,5) = 6 ways

Total number of ways = 18 + 6 = 24 ways

Required probability = 24/36 = 2/3